1. after striking the floor ,a ball rebounds 80% of the height from which it has fallen. find the total distance that it travels before coming to rest if it is dropped from a height of 120 cm.

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sougata

Total distance travel by ball is

S = 120 + 2*120*.8 + 2*120*(.8)2  + 2*120(.8)3 +.................∞

i have multiplied by 2 because ball travells two time first when ball move from flour to vertically upward and second when it come back  to flour.

S = 120 + 2*120*.8[ 1 + .8 + (.8)2+......∞]  in bracket we have infinite geometric progression

  =120 + 192[ 1/(1-0.8)]

  =120 + 192[5]

  =1080   cm


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin